OpenStudy (anonymous):

What 2 numbers multiply to equal 30 but add to be -3?

OpenStudy (anonymous):

I can't think of any.

OpenStudy (anonymous):

if your two numbers are x and y, then xy = 30 and x + y = -3. with the first equation you can get x = 30/y and subsitute x in the second equation the second equation is now 30/y + y = 3 multiply both side by y to get 30 + y² = 3y, use the quadratic formula to find that y is a complex number